OpenStudy (anonymous):

1. If you exert a force of 10.0 N to lift a box a distance of 0.75 m, how much work do you do? (1 point)0.075 J 7.5 J 10.75 J 75 J 2. If you exert a force of 500 N to walk 4 m up a flight of stairs in 4 s, how much power do you use? (1 point)31 W 500 W 2,000 W 8,000 W 3. How can you make the work output of a machine greater than the work input? (1 point)by decreasing friction by increasing the input force by increasing the output distance none of the above 4. A 100-m long ski lift carries skiers from a station at the foot of a slope to a second station 40 m abov

OpenStudy (anonymous):

I'm sorry but I dont think I can help you do this. This looks too much like an exam to me. I'd be willing to help if you gave your thoughts about the questions. Like #3, How could you get Work to increase when you put it through a machine? Cheers

OpenStudy (anonymous):

1. 10.0 N * 0.75 m = 7.5 J 2. 500 N * 4 m / 4 s = 500 W

OpenStudy (ghazi):

remember, work done in lifting an object is stored in the form of potential energy so potential energy = m*g*h = 10*0.75 = 7.5 j

OpenStudy (ghazi):

power = work done per unit time p= w/t you have 'w' you have 't' , find p

OpenStudy (ghazi):

output can never be greater than input, efficiency can be increased but output not, under ideal conditions output= input but can never be greater